OpenStudy (anonymous):

integrate of dy/sec*2y

OpenStudy (kainui):

\[\frac{1}{\sec^2y}=\cos^2y=\frac{1}{2}(1+\cos(2y))\]

OpenStudy (anonymous):

it is 1/sec^2 y

OpenStudy (kainui):

Exactly.